Daston Company manufactures two products, Product F and Product G. The company expects to produce and sell 1,630 units of Product F and 2,060 units of Product G during the current year. Data relating to the company’s three activity cost pools are given below for the current year: Total Activity Activity Cost Pools Total Cost Product F Product G Total Machine setups $ 37,290 144 setups 195 setups 339 setups Purchase orders $ 165,240 890 orders 1,150 orders 2,040 orders Order size $ 123,690 2,440 hours 4,070 hours 6,510 hours Required: Using the activity-based costing approach, determine the overhead cost per unit for each product. (Round your answers to 2 decimal places.)
Answer #1

Calculate overhead cost per unit

Product F Product G
Machine setups 37290/339*144 = 15840 37290/339*195 = 21450
Purchase orders 165240/2040*890 = 72090 165240/2040*1150 = 93150
Order size 123690/6510*2440 = 46360 123690/6510*4070 = 77330
Total overhead 134290 191930
Units 1630 2060
Overhead cost per unit 82.39 93.17